OpenStudy (anonymous):

HELPP!!! You mother and father took turns driving on a 640-mile trip that took 11 hours to complete. Your mother drove at a constant speed of 55 mph and your father drove at a constant speed of 60 mph. a) How long did your mother drove (i tried doing cross multiply and divide but that didnt work ?)

OpenStudy (anonymous):

??

OpenStudy (anonymous):

m- Hours mother drove the car and f - no. of hours father drove the car. Then acc. to ques. m+f = 11 and 55*m + 60*f = 640 (Total distance traveled) 55*m + 60*(11-m) = 640 55m + 660 - 60m = 640 5m= 660-640 5m = 20 m = 4. I hope it helps !!

OpenStudy (anonymous):

thank you